OpenStudy (destiny.duh):

Which options could be the next step in simplifying this expression? (18 + 23) + 7 Choose exactly two answers that are correct. A. 30 + 25 (use the commutative property) B. 41 + 7 (use order of operations) C. 25 + 30 (use the distributive property) D. 18 + 30 (use the associative property

OpenStudy (destiny.duh):

I will give a medal

OpenStudy (anonymous):

if you are the one who is going to solve it, how will you solve it? Remember PEMDAS

OpenStudy (destiny.duh):

48 @Data_LG2

OpenStudy (anonymous):

yes, but the question is HOW did you get 48? :) what did you do first? did you evaluate the terms inside the brackets first?

OpenStudy (destiny.duh):

yes @Data_LG2

OpenStudy (anonymous):

so if you evaluate the bracket only, what did you get?

OpenStudy (destiny.duh):

41

OpenStudy (anonymous):

so it will be 41 + 7... so which of the choices is right?

OpenStudy (destiny.duh):

b

OpenStudy (anonymous):

:)

OpenStudy (destiny.duh):

but I have to have 2 answers' @Data_LG2

OpenStudy (anonymous):

you can also apply this property of addition: \(\sf (A+B) + C =A + (B +C)\) so for your question, it will look like this: \(\sf (18+23) + 7 =18 + (23 +7)\) then do the same thing.. evaluate the bracket first for \(\sf 18 + (23 +7)\)

OpenStudy (destiny.duh):

thank u so much

OpenStudy (anonymous):

No problem ^_^ see you around!